How to i solve this equation x^3-20=44
Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

How to i solve this equation x^3-20=44
-------
x^3-64 = 0
This is a "difference of cubes" and factors as follows:
(x-4)(x^2+4x+16)=0
x=4
Solve x^2+4x+16 = 0
x = [-4 +- sqrt(16-4*16)]/2
x = [-4 +- sqrt(82)]/2
x = [-4 + sqrt(82)]/2 or x=[-4 - sqrt(82)]/2 or x=4
========================
Cheers,
Stan H.

Answer by scott8148(6628)

subtracting 44 __ x^3-64=0

this is the difference of two perfect cubes
__ it factors to (x-4)(x^2+4x+16)=0

x-4=0, so x=4

x^2+4x+16=0 does not factor, but you can find the roots using the quadratic formula